OpenStudy (bloomlocke367):

Is anybody good with conic sections? I'm kinda lost and need some help...

OpenStudy (anonymous):

Just ask your question.

OpenStudy (igreen):

^^^^^^^^^^^^^^

OpenStudy (bloomlocke367):

write an equation for the translation of x^2+y^2=49 by 3 units left and 4 units up.

OpenStudy (bloomlocke367):

@iGreen @greenlegodude57

OpenStudy (bloomlocke367):

I've narrowed it down to 2 options..

OpenStudy (igreen):

What are your options?

OpenStudy (bloomlocke367):

well all of them are (x+3)^2+(y+4)^2=49 (x+3)^2+(y-4)^2=49 (x-3)^2+(y-4)^2=49 (x-3)^2+(y+4)^2=49

OpenStudy (bloomlocke367):

I'm pretty sure it's either the first one or the second one.

OpenStudy (igreen):

Adding to 'x' will shift the graph to the left. Subtracting to 'x' will shift the graph to the right. Adding to 'y' will shift the graph down. Subtracting to 'y' will shift the graph up.

OpenStudy (igreen):

So which one do you think is correct?

OpenStudy (bloomlocke367):

the second one?

OpenStudy (bloomlocke367):

thank you. will you be able to help me more? @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

Sure, just close this and open a new one.

OpenStudy (bloomlocke367):

of course :)